OpenStudy (anonymous):

find the y intercept and the x intercept for the equation 3x- 6y = 90 whats y intercept? whats x intercept? type an ordered pair

OpenStudy (anonymous):

-6y = 90 - 3x 6y = 3x - 90 y = 3x/6 - 15 3x/6 - 15 = 0 x = 30 (x intercept) y = 3(0)/6 - 15 y = -15 (y intercept)

OpenStudy (anonymous):

thanks so what is the ordered pair

OpenStudy (anonymous):

umm the ordered pair of the x intercept is (30,0) and the one for the y is (0,-15)